OpenStudy (anonymous):

log(lower 3)((x^2)+8)=2 need the solution set

OpenStudy (anonymous):

SO since ur base for your log is 3, and it equals to two, then, 3^2 = x^2 +8 which means that x^2 + 8 = 9, where obviously x = 1 :)

OpenStudy (anonymous):

1 or -1 btw.

OpenStudy (anonymous):

because either one squared will equal 1.

OpenStudy (anonymous):

so my answer is {1,-1}